tickets for homecoming dance cost $20 for a single ticket or $35 for a couple. ticket sales totaled $2280, and 128 people attende. how many tickets of each type were sold?
Answer by Edwin McCravy(20086)

tickets for homecoming dance cost $20 for a single 
ticket or $35 for a couple. ticket sales totaled 
$2280, and 128 people attende. how many tickets of 
each type were sold?


Let x = the number of singles who attended
Let y = the number of couples who attended

Therefore the number of people who attended = 
The number of singles + twice the number of couples 
or

x + 2y people

>>...128 people attende...<<

So 

x + 2y = 128

>>...$20 for a single ticket...<<

So the money taken in from the x single tickets at
$20 a ticket was 

20x dollars.

>>...$35 for a couple...<<

So the money taken in from the y couple tickets at
$35 a ticket was

35y dollars.

>>...ticket sales totaled $2280...<<

So 20x dollars + 35y dollars = 2280 dollars or

   20x + 35y = 2280

So now you have this system of two equations in
two unknowns:


   20x + 35y = 2280
     x +  2y = 128
 
Can you solve that system?  If not post again

Answer x = 16, y = 56

16 singles and 56 couples

Edwin